OpenStudy (anonymous):

Find an equation in standard form for the ellipse with the vertical major axis of length 6, and minor axis of length 4.

OpenStudy (anonymous):

x^2/2 + y ^2/3 = 1 x^2/9 + y^2/4 = 1 x^2/3 + y^2/2 = 1 x^2/4 + y^2/9 = 1

OpenStudy (jdoe0001):

do you know what the equation of an ellipse look like?

OpenStudy (anonymous):

no

OpenStudy (jdoe0001):

standard form means, the center of it is the ORIGIN, that is, ( 0 , 0) so \(\bf \cfrac{(x-0)^2}{a^2}+\cfrac{(y-0)^2}{b^2}=1\implies \cfrac{x^2}{a^2}+\cfrac{y^2}{b^2}=1\)

OpenStudy (jdoe0001):

vertical major axis, means its LONGER over the y-axis, so that means that the "y fraction" will keep the BIGGER denominator

OpenStudy (jdoe0001):

so it'll look like 

OpenStudy (jdoe0001):

so I'd think by now you can see what the equation will look like